OpenStudy (anonymous):

For y=sin(2x+1) with 0 is less or equal to x and 180 degrees is more or equal to x how do you find the the highest and lowest points :) Thanks

OpenStudy (mathmale):

How you'd do this depends upon the approach you want to take. Are you taking a calculus course? or is this problem from trig? I'd paraphrase your question to read: "Find the maximum and minimum points of y - sin (2x + 1) on the interval\[0 \le x \le180 ~(degrees)\]

OpenStudy (mathmale):

One non-calculus approach would be to graph this function. As a general rule, if we are given y = a*sin (b*x + c), 1) The period of your sine function is 2Pi/b 2) The phase shift (horizontal shift in the basic sine graph) is -c/b 3) The amplitude of your graph is |a| (the absolute value of a). What are your a, b and c in this particular problem? Can you graph y = sin (2x+1)? If you're in calculus, find the derivative of the given sine function, and set the derivative equal to zero. Solve for the x-values that lie between 0 and 180 degrees.

OpenStudy (mathmale):

Write back with info on what you believe you still need to learn to be able to solve this problem.